# Achieving designed texture and flows in bulk active nematics using optimal control theory

### v1

@@ IV.2 Coherent flow through a ‘channel’ without confinement
With the next target state, we demonstrate controllability of the velocity field $\mathbf{u}$ of the passive solvent, through active stresses generated by extensile active nematics. We compute an activity profile that first establishes and then maintains flow of a prescribed magnitude along the $\hat{x}$ direction through a ‘channel’, that is a strip through the center of the simulation domain. To allow as much freedom as possible for the solution, the velocity is unconstrained outside the channel and there are no constraints imposed on the director anywhere. Specifically, we set $W_{\text{u}}=80$ in the channel region and zero elsewhere, and $W_{\text{Q}}=0$ everywhere. Since we have just demonstrated the ability to flatten the director field, we set the initial condition to a uniformly aligned nematic with zero velocity. We set the baseline activity to $\alpha_{0}=0.1$ so it is below the threshold for turbulence ($\alpha^{*}\approx 0.17$). We find that achieving this solution takes the controller longer than flattening the director field, and thus we set $t_{\text{F}}=200$.
Fig. 4a shows snapshots of the x-component of the velocity $u_{x}$ and the computed activity field for a trajectory integrated with the computed control solution for the low activity case. To evaluate convergence, Figs. 4a,b show the velocity and activity, each averaged over both the channel and bulk regions, as a function of time. The cost function residuals are shown as a function of time in SI Fig. 8b.
@@
The control solution drives the intrinsic active nematic bend instability [85, 86, 84], where a small bend fluctuation in the director field is reinforced by local shear, resulting in a global bend configuration and coherent flow. This configuration is closely analogous to the well-known configuration of coherent flow in a channel [77, 87], but here it is stabilized by the applied activity field instead of boundary anchoring. In particular, we see that the solution imposes uniform activity along the long axis of the subdomain $\hat{x}$, but spatial variation along the orthogonal $\hat{y}$ direction that drives and then stabilizes the bend configuration. Notice that the activity is focused in the lower half of the subdomain, but stabilizes a bend configuration and hence flow along the center of the subdomain. The target state also resembles banded states and vortex lattices that are observed in systems with significant substrate friction [88, 89, 90, 91]. However, those states are highly symmetric, with the width and number of bands determined by confinement and defect spacing. In contrast, our control solution determines the width and number of bands (in this case one) independent of the natural defect spacing.
@@
Figure 4: Coherent flow through a ‘channel’ without confining boundaries. The target state has uniform flow along the $\hat{x}$ direction with magnitude $U_{\text{T}}=0.5$ within the strip enclosed by green dashed lines. (a) The velocity field (arrows) and magnitude $u$ (color map) are shown at indicated time points for a representative trajectory under the optimal control protocol. (b) The activity field $\alpha$ (color map) and nematic director field (lines) are shown at the same time points. The initial state has the nematic uniformly aligned along the $\pi/4$ axis. (c) The flow speed $u_{x}$ (c) and the activity (d) averaged over the channel and bulk regions as a function of time. The parameter values for this solution were $\{J_{1},J_{2},J_{3},\zeta,W_{\text{u}},W_{\text{Q}},\alpha_{0},t_{\text{F}}\}=\{1.0,1.0,0.1,0.1,80.0,0,0.1,200.0\}$.
@@ IV.3 Couette flow in a channel
We further demonstrate the controllabilty of velocity field by computing an activity field that causes the system to attain a Couette flow structure in a channel that has no moving boundary. We consider a channel of width $h$, with periodic boundary conditions along the $\hat{x}$ direction, a rigid boundary at the lower wall $y=0$, and a free boundary at the upper wall $y=h$. To mimic Couette flow, we we set our target velocity profile to $\mathbf{u}^{\text{T}}_{x}=U_{\text{T}}y/h$, so that the x-component linearly increases from zero at the lower wall to a maximum $U_{\text{T}}$ at the upper wall. We impose no constraints on $u_{y}$ or $\mathbf{Q}$. Other boundary conditions are as follows. All components of $\mathbf{u}$ and $\mathbf{Q}$ are periodic in the $\hat{x}$ direction. For the lower channel wall ($y=0$), we impose no-slip conditions on the velocity, $u_{x}=u_{y}=0$. For the upper wall ($y=h$) we set $u_{y}=0$, but only no-flux (Neumann) conditions on the x-component, $\partial_{y}u_{x}=0$. We set Neumann conditions on $\mathbf{Q}$ at both walls. The dimensions of the simulation box are $60\times h$ with $h=12$. Finally, our initial condition is a uniformly aligned nematic and zero flow velocity.
@@
Figs. 5a,b show the velocity and activity fields at three time points over the course of the solution trajectory. Along the channel we find that the activity remains constant, whereas across the channel the activity profile can be computed analytically by solving the Stokes equation 2 with the assumption that Couette flow is established and the director field is initialized at $\theta=\pi/4$. Neglecting pressure gradients, the Stokes equation simplifies to, $\xi u_{x}(x,y)=-\partial_{y}(\alpha(x,y)Q_{xy})$ and $\partial_{x}\alpha(x,y)=0$, which can be simplified as $\alpha(x,y)\sim\tilde{\alpha}_{0}-\xi U_{\text{T}}y^{2}/h$, with maximum activity at $y=0$ and minimum at $y=h$ (Fig. 5c), with $\tilde{\alpha_{0}}>0$. This activity profile mimics the diffusion of momentum from a boundary that would stabilize Couette flow in the canonical moving boundary system. Interestingly, the velocity magnitude deviates below the target near the upper wall. We believe that this residual reflects the fact that a completely linear profile is incompatible with the geometry and boundary conditions that we have imposed. We found that the residual was insensitive to changing controller parameters, such as the relative magnitudes of $J_{1},J_{2},$ and $W_{\text{u}}$. Finally, we note that we obtain stable solutions only for values of $U_{\text{T}}<1.0$; higher values give rise to nucleation of defects which can disrupt the steady flow structure and thus impede convergence.
Figure 5: Couette flow in a channel without a moving boundary. (a) The velocity field (arrows) and magnitude $u$ (color map) are shown for three time points. (b) The control solution activity field $\alpha$ (color map) and nematic director field (lines) for the same time points.(c) Plot of $u_{x}$ (blue line, left y-axis) compared to the target profile (black dashed line, left y-axis), and the activity (red line, right y-axis) as a function of channel height at the final time point. The parameter values for this solution were $\{J_{1},J_{2},J_{3},\zeta,W_{\text{u}},W_{\text{Q}},t_{\text{F}}\}=\{8.0,1.0,0.1,0.1,15.0,0,15.0\}$.
@@ V Discussion and conclusions
The optimal control protocols can be implemented in any experimental system that allows external actuation of the activity (or other system parameters). For example, researchers have recently constructed light-activated microtubule-based nematics using optogenetic kinesins molecular motors [48, 49, 50, 52]. When light is shown on these motors, they bind into clusters capable of driving relative displacements of microtubules that power the active nematic; in the absence of light they do not bind and hence are inactive. Lemma et al. and Zarei et al. [48, 49] showed that the magnitude of activity is proportional to applied light intensity, and used digital light projectors to apply spatiotemporal sequences of light. Thus, the optimal control solutions obtained through our framework can be directly implemented in the light-activated nematic system as an open-loop feed-forward control protocol. Importantly, if we map the nondimensional length scale of our equations onto the light-activated nematic system (by matching the mean distance between defects), we find that our results can be obtained with applied activity fields that have spatial resolutions on the order of the experimental capability. In particular, the current experimental spatial resolution on the applied light is 70 $\upmu$M, compared to a typical defect spacing of 250 $\upmu$M [49]. In our current implementation, the spatial resolution of applied activity is the same as the spatial grade size for numerical integration, and thus we are limited to a spatial resolution of 2 non-dimensional length units in order to achieve convergence. In comparison, the typical defect spacing is 12 non-dimensional length units. However, the applied activity field can be interpolated to lower resolution without affecting the behavior of the solution, demonstrating that the experimental resolution is sufficient. It would be straightforward to impose activity on a different resolution than the numerical integration.
@@
Our previous work has shown that optimal control solutions are robust to noise, and thus can accommodate inaccuracies in the hydrodynamic theory of the active nematic system and experimental noise. If these sources of error are too large, it is straightforward to correct optimal control solutions by using feedback control. For example, one can periodically observe the state of the system along the course of the trajectory, and then either recompute the optimal control solution from the current state of the system or add a linear restoring force that drives the system to return to the computed trajectory [92].
@@
Although we have focused on active nematics in this work, the optimal control framework described here can be easily adapted to any system that can be externally actuated and for which a continuum model description can be obtained. We consider actuation of the activity field since that can be directly implemented in the light-activated nematic system [48, 49, 50, 52], but the objective function can be defined in terms of any variable that can be externally actuated. For systems that do not already have continuum model descriptions, recent work has demonstrated success in using data-driven model discovery tools to derive PDE descriptions of active matter [93, 94, 95] and other systems [96, 97, 98]. Furthermore, for systems not amenable to continuum theory, we have recently shown that optimal control solutions can be efficiently calculated for particle-based simulation models by constructing Markov state models [99, 100]. Thus, the optimal control framework can be applied to a wide variety of active and other soft material systems, both to achieve complex, functional states and to gain physical insights into the mechanisms that drive and stabilize such emergent dynamics.
@@ Acknowledgements.
This work was primarily supported by the Department of Energy (DOE) DE-SC0022291. Preliminary simulations of the system without control were supported by the National Science Foundation (NSF) DMR-1855914 and the Brandeis Center for Bioinspired Soft Materials, an NSF MRSEC (DMR-2011846). Computing resources were provided by the NSF XSEDE allocation TG-MCB090163 (Stampede and Expanse); the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award BES-ERCAP0026774; and the Brandeis HPCC which is partially supported by the NSF through DMR-MRSEC 2011846 and OAC-1920147. AB acknowledges support from NSF - 2202353 and the hospitality of the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452.

### v2

@@ IV.2 Coherent flow through a ‘channel’ without confinement
With the next target state, we demonstrate controllability of the velocity field $\mathbf{u}$ of the passive solvent, through active stresses generated by extensile active nematics. We consider a simulation box with periodic boundary conditions in both directions, with sizes $L_{x},L_{y}$. We then compute an activity profile that first establishes and then maintains flow of a prescribed magnitude along the $\hat{x}$ direction through a ‘channel’, that is a strip through the center of the simulation domain. To ensure that the result is not dependent on any commensurability between preferred wavelengths of bend deformations required to achieve flow and the periodic boundary conditions in the $\hat{y}$ direction, we performed simulations for two system sizes, with $L_{x}=100$ in both cases, and $L_{y}=50$ and $100$, respectively.
@@
To allow as much freedom as possible for the solution, the velocity is unconstrained outside the channel and there are no constraints imposed on the director anywhere. Specifically, we set $W_{\text{u}}=80$ in the channel region and zero elsewhere, and $W_{\text{Q}}=0$ everywhere. Since we have just demonstrated the ability to flatten the director field, we set the initial condition to a uniformly aligned nematic with zero velocity. We set the baseline activity to $\alpha_{0}=0.1$ so it is below the threshold for turbulence ($\alpha^{*}\approx 0.17$). We find that achieving this solution takes the controller longer than flattening the director field, and thus we set $t_{\text{F}}=200$.
Figs. 4a,b and 4e,f show snapshots of the x-component of the velocity $\mathbf{u}_{\text{x}}$ and the computed activity field for trajectories integrated with the computed control solutions for the systems with $L_{y}=50$ and $L_{y}=100$, respectively. To further show how these quantities vary across the channel direction, Figs. 4c,g show the flow speed $\langle\mathbf{u}_{\text{x}}\rangle_{x,t}$ and activity $\langle\alpha\rangle_{x,t}$ averaged over the channel axis ($\hat{x}$ direction) and time for the small and large system sizes respectively. This plot emphasizes that the activity peaks at the lower channel boundary and linearly decreases across the extent of the channel, which drives a peak in the flow velocity at the channel center. To evaluate convergence, Figs. 4d,h show the flow speed averaged over the channel region and exterior region as a function of time. We see that the fluid in the exterior becomes nearly quiescent at very early times, but the flow speed in the channel does not saturate until $t>100$. To present more information about convergence, the cost function residuals are shown as a function of time in SI Fig. 10b.
@@
The control solution drives the intrinsic active nematic bend instability Thampi, Golestanian, and Yeomans (2014); Giomi et al. (2014); Marchetti et al. (2013), where a small bend fluctuation in the director field is reinforced by local shear, resulting in a global bend configuration and coherent flow. This configuration is closely analogous to the well-known configuration of coherent flow in a channel Wagner et al. (2022); Shendruk et al. (2017), but here it is stabilized by the applied activity field instead of boundary anchoring. In particular, we see that the solution imposes uniform activity along the long axis of the subdomain $\hat{x}$, but spatial variation along the orthogonal $\hat{y}$ direction that drives and then stabilizes the bend configuration. Notice that for both system sizes the activity is focused in the lower half of the subdomain, but stabilizes a bend configuration and hence flow along the center of the subdomain. Also, the bend deformation has a similar wavelength in both system sizes, commensurate with the channel width, but the larger system size adopts a large region of uniform nematic below the channel.
The target states also resemble banded states and vortex lattices that are observed in systems with significant substrate friction Caballero, You, and Marchetti (2023); Doostmohammadi et al. (2018); Doostmohammadi, Thampi, and Yeomans (2016); Schimming, Reichhardt, and Reichhardt (2024b). However, those states are highly symmetric, with the width and number of bands determined by confinement and defect spacing. In contrast, our control solution determines the width and number of bands (in this case one) independent of the natural defect spacing.
@@
Figure 6: Robustness of the control solution to adding noise to the initial condition. (a) Deviation as a function of time between the trajectories integrated computed with and without noise ($\mathcal{J}_{\text{T}}$, Eq. (9)). We added Gaussian noise to the initial condition with indicated magnitude $\epsilon_{\text{n}}$, and integrated the dynamics using the control protocol computed in the absence of noise ($\epsilon_{\text{n}}=0$) that is shown in Fig. 4. The inset shows the deviation of the final state $\mathcal{J}_{\text{T}}(t_{\text{F}})$ as a function of noise magnitude $\epsilon_{\text{n}}$. (b,c) The final states for (b) small ($\epsilon_{\text{n}}=0.05$) and (c) large ($\epsilon_{\text{n}}=0.30$) noise. We see robust flow patterns even when the error is large in the end state for 30% noise.
@@ IV.3 Couette flow in a channel
We further demonstrate the controllabilty of velocity field by computing an activity field that causes the system to attain a Couette flow structure in a channel that has no moving boundary. We consider a channel of width $h$, with periodic boundary conditions along the $\hat{x}$ direction, a rigid boundary at the lower wall $y=0$, and a free boundary at the upper wall $y=h$. To mimic Couette flow, we set our target velocity profile to $\mathbf{u}^{\text{T}}_{x}=U_{\text{T}}y/h$, so that the x-component linearly increases from zero at the lower wall to a maximum $U_{\text{T}}$ at the upper wall. We impose no constraints on $u_{y}$ or $\mathbf{Q}$. Other boundary conditions are as follows. All components of $\mathbf{u}$ and $\mathbf{Q}$ are periodic in the $\hat{x}$ direction. For the lower channel wall ($y=0$), we impose no-slip conditions on the velocity, $u_{x}=u_{y}=0$. For the upper wall ($y=h$) we set $u_{y}=0$, but only no-flux (Neumann) conditions on the x-component, $\partial_{y}u_{x}=0$. We set Neumann conditions on $\mathbf{Q}$ at both walls. The dimensions of the simulation box are $60\times h$ with $h=12$. Finally, our initial condition is a uniformly aligned nematic and zero flow velocity. Figs. 5a,b show the velocity and activity fields at three time points over the course of the solution trajectory. Along the channel we find that the activity remains constant, whereas across the channel the activity profile can be computed analytically by solving the Stokes equation 2 with the assumption that Couette flow is established and the director field is initialized at $\theta=\pi/4$. Neglecting pressure gradients, the Stokes equation simplifies to, $\xi u_{x}(x,y)=-\partial_{y}(\alpha(x,y)Q_{xy})$ and $\partial_{x}\alpha(x,y)=0$, which can be simplified as $\alpha(x,y)\sim\tilde{\alpha}_{0}-\xi U_{\text{T}}y^{2}/h$, with maximum activity at $y=0$ and minimum at $y=h$ (Fig. 5c), with $\tilde{\alpha_{0}}>0$. This activity profile mimics the diffusion of momentum from a boundary that would stabilize Couette flow in the canonical moving boundary system. Interestingly, the velocity magnitude deviates below the target near the upper wall. We believe that this residual reflects the fact that a completely linear profile is incompatible with the geometry and boundary conditions that we have imposed. We found that the residual was insensitive to changing controller parameters, such as the relative magnitudes of $J_{1},J_{2},$ and $W_{\text{u}}$. Finally, we note that we obtain stable solutions only for values of $U_{\text{T}}<1.0$; higher values give rise to nucleation of defects which can disrupt the steady flow structure and thus impede convergence.
Figure 7: Robustness of the control solution to the coarseness of control field (a) Deviation as a function of time between the trajectories integrated computed with coarser and optimal $\alpha$ ($\mathcal{J}_{\text{T}}$, Eq. (9)). Fig. 4. The inset shows the deviation of the final state $\mathcal{J}_{\text{T}}(t_{\text{F}})$ as a function $\delta\alpha$. (b,c) The initial and final states and activity profile for $\delta_{\alpha}=4$. We see low errors for $\delta_{\alpha}\lesssim 12$, which is the defect spacing at these parameters.
IV.4 Robustness of the optimal control solution to perturbations.
While the optimal control solution in this work is computed for a deterministic PDE model, the ultimate goal is to apply that solution to an experimental system. Thus, it is important to understand how robust the computed solution is to perturbations, since models are not completely accurate and experiments inevitably have noise and measurement errors.
In previous work Ghosh et al. (2024), we found that the optimal control solution to an active polar fluid was robust to noise. We observe the same robustness in this active nematic system. In particular, we focus on the coherent flow solution shown in Fig. 4. We perturb the initial condition by adding Gaussian noise with magnitude $\epsilon_{\text{n}}$ to the values of $\mathbf{Q}$ and $\mathbf{u}$ at each lattice site, and then integrate the forward solution of the dynamics with the optimal control solution that we computed without noise. As a measure of error, Fig. 6 shows the deviation of the trajectory in the presence of noise compared to the original trajectory without noise 4
$\mathcal{J}_{\text{T}}(t)=$ $\frac{1}{\Omega}\int_{\Omega}d\mathbf{r}\frac{1}{2}\left[(\mathbf{Q}(\mathbf{r},t)-\mathbf{Q}^{*}(\mathbf{r},t))^{2}\right.+$
$\left.(\mathbf{u}(\mathbf{r},t)-\mathbf{u}^{*}(\mathbf{r},t))^{2}\right],$ (9)
where $(\mathbf{Q}^{*}(\mathbf{r},t),\mathbf{u}^{*}(\mathbf{r},t))$ is the trajectory in Fig. 4 and the error is normalized by the spatial extent $\Omega$. The observation that we obtain consistent trajectory even for large perturbations to the initial condition suggests that the optimal control solution is robust to noise. However, we consider approaches for cases in which the solution becomes inaccurate due to errors in section V. Importantly, at all noise values deviations remain extremely small until about the middle of the trajectory, $t=100$, indicating that the error accumulates over time for it affects the solution. This trend suggests that the feedback approaches to address noise or errors discussed in section V can be applied successfully.
We also investigated the sensitivity of the control solution to the spatial resolution of the applied light field, since only limited resolution will be available in experimental systems (see section V for a discussion of the available resolution in the experimental light activated nematic system). In this case, we use the flattening of the domain solution shown in Fig. 7. We coarse-grain the optimal activity $\alpha$ onto a coarser grid with ${L_{x}}/{\delta_{\alpha}}\times{L_{y}}/{\delta_{\alpha}}$ grid points, where $\delta_{\alpha}$ is the coarse-graining parameter. The lower spatial resolution $\alpha$ is smoothed by applying a multidimensional Gaussian filter. The filter computes a weighted sum of neighboring points based on a Gaussian kernel, where the weights are determined by the Gaussian function $G(x)={1}/{\sqrt{2\pi\sigma^{2}}}e^{-\frac{x^{2}}{2\sigma^{2}}}$. The Gaussian filter is applied independently along each axis, with the 2D filter expressed as: $\alpha_{\delta_{\alpha}}=\sum_{i=-K}^{K}\sum_{j=-K}^{K}G(i,j)\alpha_{\text{coarse}}$ where $G(i,j)$ is the Gaussian kernel. The parameter $\sigma=0.7$ and kernel size $K=4$ controls the spread of the filter Virtanen et al. (2020). We then integrate the forward solution using the interpolated activity field. Fig. 7 shows the deviation from the trajectory computed with the original resolution (Eq. (9)) as a function of $\delta_{\alpha}$. We see that deviations from the original trajectory remain relatively small until the resolution becomes on the order of the defect spacing, $\delta_{\alpha}\approx 12$. This threshold makes sense, as it sets the characteristic length-scale for gradients in the nematic texture that control collective defect motions (as opposed to the shorter-scale gradients at the core of a defect). Moreover, the solutions still achieve the objective (uniform nematic in the region of interest) reasonably well (Fig. 7b, c).
Finally, if one knows the experimental limit on resolution of the applied field a priori, one can compute the optimal control solution with the activity field constrained to that resolution. That currently is not possible in our implementation, since the resolution of the activity field is fixed to the finite element grid, but it would be straightforward to include this feature in the framework.
@@ V Discussion and conclusions
The optimal control protocols can be implemented in any experimental system that allows external actuation of the activity (or other system parameters). For example, researchers have recently constructed light-activated microtubule-based nematics using optogenetic kinesins molecular motors Lemma et al. (2023); Zarei et al. (2023); Zhang et al. (2021); Ross et al. (2019). When light is shone on these motors, they bind into clusters capable of driving relative displacements of microtubules that power the active nematic; in the absence of light they do not bind and hence are inactive. Lemma et al. and Zarei et al. Lemma et al. (2023); Zarei et al. (2023) showed that the magnitude of activity is proportional to applied light intensity, and used digital light projectors to apply spatiotemporal sequences of light. Thus, the optimal control solutions obtained through our framework can be directly implemented in the light-activated nematic system as an open-loop feed-forward control protocol. Importantly, if we map the nondimensional length scale of our equations onto the light-activated nematic system (by matching the mean distance between defects), we find that our results can be obtained with applied activity fields that have spatial resolutions on the order of the experimental capability. In particular, the current experimental spatial resolution on the applied light is 70 $\upmu$m, compared to a typical defect spacing of 250 $\upmu$mZarei et al. (2023). As shown in Fig. 7, the computed activity field can be interpolated to coarser scales as large as the typical defect spacing ($\approx 12$ non-dimensional length units), suggesting that the experimental resolution is more than sufficient.
@@
Our previous work has shown that optimal control solutions are robust to noise, and thus can accommodate inaccuracies in the hydrodynamic theory of the active nematic system and experimental noise. If these sources of error are too large, it is straightforward to correct optimal control solutions by using feedback control. For example, one can periodically observe the state of the system along the course of the trajectory, and then either recompute the optimal control solution from the current state of the system or add a linear restoring force that drives the system to return to the computed trajectory Bechhoefer (2021). As shown in Fig. 6, errors do not accumulate significantly until timescales on order of half the trajectory, suggesting that feedback would not have to be implemented too frequently.
@@
Although we have focused on active nematics in this work, the optimal control framework described here can be easily adapted to any system that can be externally actuated and for which a continuum model description can be obtained. We consider actuation of the activity field since that can be directly implemented in the light-activated nematic system Lemma et al. (2023); Zarei et al. (2023); Zhang et al. (2021); Ross et al. (2019), but the objective function can be defined in terms of any variable that can be externally actuated. For systems that do not already have continuum model descriptions, recent work has demonstrated success in using data-driven model discovery tools to derive PDE descriptions of active matter Joshi et al. (2022); Supekar et al. (2023); Golden et al. (2023) and other systems Fuhg, Jones, and Bouklas (2024); Linden, Kramer, and Rangamani (2022); Ramezanian-Panahi et al. (2022). Furthermore, for systems not amenable to continuum theory, we have recently shown that optimal control solutions can be efficiently calculated for particle-based simulation models by constructing Markov state models Trubiano and Hagan (2022, 2024). Thus, the optimal control framework can be applied to a wide variety of active and other soft material systems, both to achieve complex, functional states and to gain physical insights into the mechanisms that drive and stabilize such emergent dynamics.
@@ Acknowledgements.
This work was primarily supported by the Department of Energy (DOE) DE-SC0022291. Preliminary simulations of the system without control were supported by the National Science Foundation (NSF) DMR-1855914 and the Brandeis Center for Bioinspired Soft Materials, an NSF MRSEC (DMR-2011846). Computing resources were provided by the NSF XSEDE allocation TG-MCB090163 (Stampede and Expanse); the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award BES-ERCAP0026774; and the Brandeis HPCC which is partially supported by the NSF through DMR-MRSEC 2011846 and OAC-1920147. AB acknowledges support from NSF - 2202353 and the hospitality of the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. SG would like thank Fernando Caballero, Michael M. Norton, Chaitanya Joshi and Suraj Shankar for helpful discussions.